OpenStudy (anonymous):

evaluate log4 (1/32)

zepdrix (zepdrix):

\[\Large \log_4\left(\frac{1}{32}\right)\]We gotta figure this out? Hmmm

zepdrix (zepdrix):

I guess we could ummm do something like this,\[\Large \log_4\left(\frac{1}{32}\right)\quad=\quad \log_4\left(\frac{1}{16}\cdot\frac{1}{2}\right)\]And then using a rule of logs:\[\Large \color{teal}{\log(a\cdot b)\quad=\quad \log(a)+\log(b)}\]We can write it as:

zepdrix (zepdrix):

\[\Large \log_4\left(\frac{1}{16}\cdot\frac{1}{2}\right)\quad=\quad \log_4\left(\frac{1}{16}\right)+\log_4\left(\frac{1}{2}\right)\]

zepdrix (zepdrix):

Should be a little bit easier to work with from here. \[\Large \frac{1}{16}\quad=\quad \frac{1}{4^2}\quad=\quad 4^{-2}\]Do you understand how we're able to write the 1/16 like this?

OpenStudy (anonymous):

yes thnx!

zepdrix (zepdrix):

The 1/2 is a little tricky :)\[\Large \frac{1}{2}\quad=\quad \frac{1}{\sqrt4}\quad=\quad \frac{1}{4^{1/2}}\quad=\quad 4^{-1/2}\]Do you understand how to solve the problem from there? :o

OpenStudy (anonymous):

actually got confused there

zepdrix (zepdrix):

heh :)

zepdrix (zepdrix):

\[\Large 2\quad=\quad \sqrt4\]Ok with that part? :o

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Don't remember how to write roots as rational expressions? :)

OpenStudy (anonymous):

kinda

zepdrix (zepdrix):